OpenStudy (anonymous):

***Answer for a medal please*** Two hockey pucks on an ice rink are held together with a compressed spring between them. The pucks are released and the spring pushes them in opposite directions. One puck of mass 0.5 kg moves at 8 m/s. Calculate the speed of the other puck of mass 2 kg. _________ m/s -2 +2 -4 +4

OpenStudy (anonymous):

@raghuvarenya???

OpenStudy (anonymous):

or somebody who is capable of answering...

OpenStudy (anonymous):

use law of conversation of Momentum; \[m_1u_1 + m_2u_2 = m_1v_1 + m_2v_2 \]

OpenStudy (anonymous):

can you help me plug the stuff in and i can solve it from there....

OpenStudy (anonymous):

where the m1 and m2 is the mass of two objects. u1 and u2 are the initial velocity of hockey pucks v1 and v2 are the final velocity of hockey pucks

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

wait im still having a hard time plugging the stuff into the equation... @MrSensitive

OpenStudy (anonymous):

try to rearrange the equation

OpenStudy (anonymous):

i did... m1=0.5 thats all i have...

OpenStudy (anonymous):

m2 = 2kg the initial velocity of both pucks are 0 the v1=8m/s and all u have to do is find v2

OpenStudy (anonymous):

do i add and subtract to find v2?

OpenStudy (anonymous):

0 = 4 - 2v -4 = -2v v= +2m/s

OpenStudy (anonymous):

great! now i solve right?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

the answer +2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but don't relay on me though

OpenStudy (anonymous):

ok thanks for the help though!